THE ROTATION-TORSION SPECTRUM^{a} OF CH_{2}DOH

A. EL HILALI, \underline{L. H. COUDERT}, LISA, CNRS/Universités Paris Est et Paris Diderot, 61 Avenue du Général de Gaulle, 94010 Créteil, France; L. MARGULÈS, R. MOTIYENKO, Laboratoire PhLAM, UMR 8523 CNRS, Bât. P5, Université des Sciences et Technologies de Lille 1, 59655 Villeneuve d’Ascq Cedex, France; and S. KLEE, Physikalisch-Chemisches Institut, Justus-Liebig-Universität Gießen, 35392 Gießen, Germany.

Due to the asymmetry of the CH_{2}D group, the internal rotation problem in the partially deuterated species of methanol CH_{2}DOH is a complicated one as, unlike in the normal species CH_{3}OH, the inertia tensor depends on the angle of internal rotation. The CH_{2}DOH species also displays a dense far infrared torsional spectrum difficult to assign. Recently 38 torsional subbands of CH_{2}DOH have been identified,^{b} but for most of them there is neither an assignment nor an analysis of their rotational structure. In this paper an analysis of the rotation-torsion spectrum of CH_{2}DOH will be presented. The rotational structure of 23 torsional subbands have been assigned. These subbands are Δv_{t} ≥ 1 perpendicular subbands with a value of v′_{t} up to 10^{b} and values of K′ and K′′ ranging from 0 to 9. For all subbands, the Q-branch was assigned, for 3 subbands, the R- and P -branches could also be found. The results of the rotational analysis with an expansion in J(J + 1) of the new subbands and of already observed ones^{c} will be presented. When available, microwave lines within the lower torsional level, recorded in this work or already measured,^{d} were added to the data set. A theoretical approach aimed at calculating the rotation-torsion energy levels has also been developed. It is based on an expansion in terms of rotation-torsion operators with C_{s} symmetry and accounts for the dependence of the inertia tensor on the angle of internal rotation. This approach will be used to carry out a preliminary global analyses of the wavenumbers and of the frequencies.

aThis work is supported by the ANR-08-BLAN-0225 contract.

bLauvergnat, Coudert, Klee, and Smirnov, J. Mol. Spec. 256 (2009) 204.

cQuade, Liu, Mukhopadhyay, and Su, J. Mol. Spec. 192 (1998) 378; and Mukhopadhyay, J. Mol. Struct. 695-696 (2004) 357.

dLiu and Quade, J. Mol. Spec. 146 (1991) 252; and Mukhopadhyay et al., J. Chem. Phys. 116 (2002) 3710.